# MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought Reasoning enhances Formal Theorem Proving

### v1

@@ Abstract
Solving mathematical problems using computer-verifiable languages like Lean has significantly impacted mathematical and computer science communities. State-of-the-art methods utilize single Large Language Models (LLMs) as agents or provers to either generate complete proof or perform tree searches. However, single-agent methods inherently lack a structured way to combine high-level reasoning in Natural Language (NL) with Formal Language (FL) verification feedback. To solve these issues, we propose MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought framework, (to the best of our knowledge), the first multi-agent framework for Lean4 theorem proving that balance high-level NL reasoning and FL verification in Long CoT. Using this structured interaction, our approach enables deeper insights and long-term coherence in proof generation, with which past methods struggle. We do this by leveraging emergent formal reasoning ability in Long CoT using our novel LoT-Transfer Learning training-inference pipeline. Extensive experiment shows that our framework achieves 54.51% accuracy rate on the Lean4 version of MiniF2F-Test dataset, largely outperforming GPT-4 (22.95%), single-agent tree search (InternLM-Step-Prover, 50.70%), and whole-proof generation (DeepSeek-Prover-v1.5, 48.36%) baselines. Furthermore, our findings highlight the potential of combining Long CoT with formal verification for a more insightful generation in a broader perspective.
@@ 1 Introduction
Extensive experiments demonstrate that MA-LoT framework effectively enhances the model’s formal reasoning ability through multi-agent design and emergent formal reasoning in Long CoT. The framework can successfully proof some of the advanced IMO, and AIME problems in the MiniF2F datset (Zheng et al., 2021), with which existing models struggle. Under similar sampling budgets, our framework achieves a 54.51% accuracy rate, surpassing state-of-the-art whole-proof generation models (DeepSeek-Prover-V1.5(Xin et al., 2024b), 48.36%) and tree-search baseline (InternLM-Step-Prover(Wu et al., 2024a), 50.70%).
@@ 3 Experiments
LoT (whole-proof) 7B pass@128 62.70% 52.05% 57.42%
MA-LoT $64+32\times 2$ 64.34% 54.51% 59.22%
MA-LoT cumulative 65.98% 56.97% 61.48%
Table 1: Main experimental results of MA-LoT. LoT (whole-proof) indicates the whole-proof generation results of our LoT-Solver model, MA-LoT with sampling budget $64\times 2$ is the major result of our MA-LoT. For baselines using tree-search baseline, we take the closest available sample budget result, for baselines using whole-proof generation, we take the pass@128 sample budget.
@@ 3.1.2 Baselines
To highlight MA-LoT ’s capabilities, we select some of the most competitive baselines in recent years, covering both tree-search and whole-proof generation approaches. For tree-search methods, we include: Expert Iteration (Polu et al., 2022), Llemma (Azerbayev et al., 2023), ReProver (Yang et al., 2024b), Lean-STaR(Lin et al., 2024), and InternLM2.5-StepProver(Wu et al., 2024a) as our baselines. For whole-proof generation baselines, we include closed-source LLMs, represented by GPT-4-Turbo(Achiam et al., 2023) and Gemini-1.5(Reid et al., 2024), and open-source expert models, such as DeepSeek-Math(Shao et al., 2024), TheoremLlama(Wang et al., 2024), and DeepSeek-Prover-v1.5-RL (Xin et al., 2024b).
For whole-proof generation baselines, we set the sample budget to pass@128 with 4096 context length, balancing robustness and manageable GPU consumption. For tree-search methods, we align the search cost as closely as possible to whole-proof generation. [^2] This explains why we exclude RMaxTS for DeepSeek-Prover-v1.5, as its smallest disclosed sample budget ($1\times 3200$) not a comparable result to our baselines.
@@ 3.2 Implementation Details
In the training process of the model, we use Openo1-SFT-Pro, LoT-ProverData, and LoT-CorrectionData to train the base model DeepSeek-Prover-v1.5-SFT. For different training stages, the learning rate is as follows: 1E-5 for NL Long CoT training, 1E-7 for LoT-TL on LoT-PD, and 1E-6 for LoT-CD. The total computational cost for training is around 1 GPU day, and evaluation is 11 GPU days on a 4×4$\times$4 × H100-96G cluster [^3] The high evaluation cost due to accelerate inference methods like vLLM unable to fit our machine. To evaluate our framework in detail, we present three sets of results, including: (1) LoT (whole-proof): pass@128 whole-proof generation result of LoT-Solver. (2) MA-LoT: Our primary evaluation result, where the prover performs 64 whole-proof generations and undergoes two rounds of corrector refinement [^4] Because we don’t need to pass correct proof to corrector, two rounds of correction is approximately the same as one round of whole-proof generation. (3) Cumulative Results: A combined evaluation aggregating all LoT models outputs obtained throughout the experiment process.
@@ 3.3 Results
Table 1 presents our main results, showing that MA-LoT achieves 54.51% accuracy rate on MiniF2F-Test benchmark, and 52.05% for LoT-Solver using whole proof-completion. Detailed analysis also shows that our models can solve some IMO and AIME problems that previous models struggled with. Both models surpass state-of-the-art tres-search (InternLM2.5) and whole-proof generation (DeepSeek-Prover-v1.5) baselines, demonstrating that our proposed multi-agent framework based on Long CoT excels in formal theorem proving.
MA-LoT outperforms all tree-search baselines by at least 3.81% because its prover agent constructs proofs with high-level NL planning using emergent Lean Long CoT reasoning capability. This indicates our prover agent can leverage LLMs’ strong NL reasoning ability, which leads to more comprehensive proofs. Additionally, MA-LoT surpasses whole-proof generation baselines by at least 6.15%, as its corrector agent analyzes, reflects, and reformulates proofs based on Lean4 executor feedback in Long CoT. The strong performance also demonstrates the effectiveness of our ideas of integrating FL verification in NL Long CoT reasoning with its emergent capability. Notably, although MA-LoT is based on DeepSeek-Prover-v1.5-SFT, it outperforms its RL-trained variant by 6.15%. This suggests that our multi-agent training framework and Lean-based Long CoT methodology align more naturally with formal theorem proving than RL alone.
@@
MiniF2F-Test 51.64% 53.28% 54.51% 55.33%
Table 2: Result of different rounds of corrector agent correction
@@
Table 3: Ablation study result in pass@64
@@ 3.4 Corrector Agent study
To assess the impact of the corrector agent in our multi-agent system, we present the cumulative accuracy on MiniF2F across different rounds in Table 2. The prover column represents the pass@64 accuracy rate of the prover agent in whole-proof generation, while Round-$i$ columns indicate successive correction rounds, each permitting up to 64 refinements per theorem. The results indicate during the first three correction rounds, the corrector agent successfully refines an average of 1.12% of theorems. Our analysis shows that most corrected proofs belong to IMO, AIME, and high-level MATH problems, which are particularly difficult for prior models. This highlights the corrector agent’s ability to analyze feedback from the Lean4 executor feedback using emergent Lean capability in Long CoT to discover non-trivial proofs. The case study for the analysis and regeneration of new proofs can be found in Section 3.6 and Appendix D.
@@ 3.5 Ablation Study
To evaluate the effectiveness of each component of our LoT-TL training pipeline, we conduct this thorough ablation study. We demonstrate that the elements in LoT-TL pipeline work synergistically to strengthen the model’s formal theorem-proving capability through the integration of FL in Long CoT. We apply the pass@64 accuracy rate on the whole-proof generation method for this set of experiments. The results are presented in Table 3.
@@ 5 Conclusion
This paper introduces MA-LoT, (to the best of our knowledge), the first multi-agent Lean-based Long Chain-of-Thought framework for formal theorem proving. Our approach addresses the limitations of single-agent systems, which either under-utilize the NL reasoning and planning capabilities of LLMs or fail to integrate formal verification feedback effectively. By structuring interactions between a prover and a corrector agent through Long CoT, MA-LoT enables deeper insights and long-term coherence in proof generation. To support this framework, we propose LoT-TL, a training pipeline that makes formal reasoning Long CoT capability emerge to LLMs without requiring annotated Lean Long CoT data. Through extensive experiments on the MiniF2F-Test benchmark, MA-LoT achieves 54.51% accuracy, surpassing all baselines, including both tree-search and whole-proof generation methods. These results underscore the advantages of integrating formal verification with structured reasoning leading to better Lean4 theorem proving capability. Beyond theorem proving, LoT-TL training pipeline demonstrates a potential method for applying Long CoT techniques to domain-specific tasks without specialized annotations. Additionally, the success of multi-agent Long CoT in Lean4 suggests broader applications of formal verification for enhancing structured reasoning across diverse fields.

### v2

@@ Abstract
Solving mathematical problems using computer-verifiable languages like Lean has significantly impacted mathematical and computer science communities. State-of-the-art methods utilize single Large Language Models (LLMs) as agents or provers to either generate complete proof or perform tree searches. However, single-agent methods inherently lack a structured way to combine high-level reasoning in Natural Language (NL) with Formal Language (FL) verification feedback. To solve these issues, we propose MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought framework, (to the best of our knowledge), the first multi-agent framework for Lean4 theorem proving that balance high-level NL reasoning and FL verification in Long CoT. Using this structured interaction, our approach enables deeper insights and long-term coherence in proof generation, with which past methods struggle. We do this by leveraging emergent formal reasoning ability in Long CoT using our novel LoT-Transfer Learning training-inference pipeline. Extensive experiment shows that our framework achieves 61.07% accuracy rate on the Lean4 version of MiniF2F-Test dataset, largely outperforming GPT-4 (22.95%), single-agent tree search (InternLM-Step-Prover, 50.70%), and whole-proof generation (Godel-Prover, 55.33%) baselines. Furthermore, our findings highlight the potential of combining Long CoT with formal verification for a more insightful generation in a broader perspective.
@@ 1 Introduction
Extensive experiments demonstrate that MA-LoT framework effectively enhances the model’s formal reasoning ability through multi-agent design and emergent formal reasoning in Long CoT. The framework can successfully proof some of the advanced IMO, and AIME problems in the MiniF2F datset (Zheng et al., 2021), with which existing models struggle. Under similar sampling budgets, our framework achieves a 61.07% accuracy rate, surpassing state-of-the-art whole-proof generation models (Godel-Prover (Lin et al., 2025), 55.33%) and tree-search baseline (InternLM-Step-Prover(Wu et al., 2024a), 50.70%).
@@ 3 Experiments
STP-Lean (Dong & Ma, 2025) 7B - 56.15% -
Godel-Prover (Lin et al., 2025) 7B pass@32 - 55.33% -
Ours
DeepSeek-Prover + LoT (whole-proof) 7B pass@128 62.70% 52.05% 57.42%
DeepSeek-Prover + MA-LoT $64+32\times 2$ 64.34% 54.51% 59.22%
Godel-Prover + LoT (whole-proof) pass@32 - 57.79% -
Godel-Prover + MA-LoT $16+8\times 2$ 61.07% -
MA-LoT cumulative 65.98% 63.93% 64.96%
Table 1: Main experimental results of MA-LoT. Our results is presented as Base model + method. The LoT (whole-proof) indicates using the whole-proof results of our LoT-Solver model and MA-LoT indicates the result of our whole pipeline. The sampling budget $x+k\times y$ in the MA-LoT framework indicates we first perform $x$ whole-proof writing using the prover agent and $k$ rounds of correction using the corrector agent. In practice, one round of correction costs approximately half of the whole-proof generation budget, result in $y=\frac{1}{2}x$.
@@ 3.1.2 Baselines
To highlight MA-LoT ’s capabilities, we select some of the most competitive baselines in recent years, covering both tree-search and whole-proof generation approaches. For tree-search methods, we include: Expert Iteration (Polu et al., 2022), Llemma (Azerbayev et al., 2023), ReProver (Yang et al., 2024b), Lean-STaR(Lin et al., 2024), and InternLM2.5-StepProver(Wu et al., 2024a) as our baselines. For whole-proof generation baselines, we include closed-source LLMs, represented by GPT-4-Turbo(Achiam et al., 2023) and Gemini-1.5(Reid et al., 2024), and open-source expert models, such as DeepSeek-Math(Shao et al., 2024), TheoremLlama(Wang et al., 2024), DeepSeek-Prover-v1.5-RL (Xin et al., 2024b), STP-Lean (Dong & Ma, 2025), and Godel-Prover-SFT (Lin et al., 2025).
For whole-proof generation baselines, we set the sample budget to pass@128 with 4096 context length (expect for Godel-Prover, where, following the original paper, we use pass@32), balancing robustness and manageable GPU consumption. For tree-search methods, we align the search cost as closely as possible to whole-proof generation. [^2] This explains why we exclude RMaxTS for DeepSeek-Prover-v1.5, as its smallest disclosed sample budget ($1\times 3200$) not a comparable result to our baselines.
@@ 3.2 Implementation Details
In the training process of the model, we use Openo1-SFT-Pro, LoT-ProverData, and LoT-CorrectionData to train two base models, namely DeepSeek-Prover-v1.5-SFT and Godel-Prover-SFT. For different training stages, the learning rate is as follows: 1E-5 for NL Long CoT training, 1E-7 for LoT-TL on LoT-PD, and 1E-6 for LoT-CD. The total computational cost for training is around 1 GPU day, and evaluation is 11 GPU days on a 4×4$\times$4 × H100-96G cluster [^3] The high evaluation cost due to accelerate inference methods like vLLM unable to fit our machine. To evaluate our framework in detail, we present three sets of results, including: (1) LoT (whole-proof): pass@128 whole-proof generation result of LoT-Solver. (2) MA-LoT: Our primary evaluation result, where the prover performs 64 whole-proof generations and undergoes two rounds of corrector refinement [^4] Because we don’t need to pass correct proof to corrector, two rounds of correction is approximately the same as one round of whole-proof generation. (3) Cumulative Results: A combined evaluation aggregating all LoT models outputs obtained throughout the experiment process.
@@ 3.3 Results
Table 1 presents our main results, showing that MA-LoT achieves 61.07% accuracy rate on MiniF2F-Test benchmark and 57.79% for LoT-Solver using whole-proof generation. After the enhancement of MA-LoT, we achieve 10.37% on Godel-Prover and 12.72% on DeepSeek-Prover on the MiniF2F-Test benchmark. This significant and uniform improvement of our framework indicates its effectiveness. Detailed analysis also shows that our models can solve IMO and AIME problems that previous models struggled with. Our model surpasses state-of-the-art tree-search (InternLM-2.5) and whole-proof generation (Godel-Prover) baselines, demonstrating that our proposed multi-agent framework based on Long CoT excels in formal theorem proving.
MA-LoT outperforms all tree-search baselines by at least 20.45% because its prover agent constructs proosfs with high-level NL planning using emergent Lean Long CoT reasoning capability. This indicates our prover agent can leverage LLMs’ strong NL reasoning ability, which leads to more comprehensive proofs. Additionally, MA-LoT surpasses whole-proof generation baselines by at least 10.37%, as its corrector agent analyzes, reflects, and reformulates proofs based on Lean4 executor feedback in Long CoT. The strong performance also demonstrates the effectiveness of our ideas of integrating FL verification in NL Long CoT reasoning with its emergent capability. Notably, the DeepSeek-Prover + MA-LoT is based on DeepSeek-Prover-v1.5-SFT; it outperforms its RL-trained variant by 6.15%. This suggests that our multi-agent training framework and Lean-based Long CoT methodology align more naturally with formal theorem proving than RL alone.
@@
DS-Prover-v1.5 51.64% 53.28% 54.51% 55.33%
Godel 54.92% 59.43% 61.07% 61.89%
Table 2: Result of different rounds of corrector agent correction with DeepSeek-Prover and Godel-Prover as base models
@@
Table 3: Ablation study result in pass@64 under DeepSeek-Prover as base model
@@ 3.4 Corrector Agent study
To assess the impact of the correct agent in MA-LoT, we present the cumulative accuracy on the MiniF2F-Test across different rounds of correction in Table 2 for two of our base models. The prover column presents the pass@64 (or 16) accuracy rate of the prover agent in whole-proof generation, while Round-$i$ columns indicate successive correction rounds. The results indicate that during the first three correction rounds, the corrector agent successfully correct an average of 11.55% of theorems that the prover unable to answer. Our analysis shows that most corrected proofs belong to IMO, AIME, and high-level MATH problems, which are particularly difficult for prior models. This highlights the corrector agent’s ability to analyze feedback from the Lean4 executor feedback using emergent Lean capability in Long CoT to discover non-trivial proofs. The case study for the analysis and regeneration of new proofs can be found in Section 3.6 and Appendix D.
@@ 3.5 Ablation Study
To evaluate the effectiveness of each component of our LoT-TL training pipeline, we conduct this thorough ablation study. We demonstrate that the elements in LoT-TL pipeline work synergistically to strengthen the model’s formal theorem-proving capability through the integration of FL in Long CoT. We use DeepSeek-Prover-v1.5-SFT as our base model and apply the pass@64 accuracy rate on the whole-proof generation method for this set of experiments. The results are presented in Table 3.
@@ 5 Conclusion
This paper introduces MA-LoT, (to the best of our knowledge), the first multi-agent Lean-based Long Chain-of-Thought framework for formal theorem proving. Our approach addresses the limitations of single-agent systems, which either under-utilize the NL reasoning and planning capabilities of LLMs or fail to integrate formal verification feedback effectively. By structuring interactions between a prover and a corrector agent through Long CoT, MA-LoT enables deeper insights and long-term coherence in proof generation. To support this framework, we propose LoT-TL, a training pipeline that makes formal reasoning Long CoT capability emerge to LLMs without requiring annotated Lean Long CoT data. Through extensive experiments on the MiniF2F-Test benchmark, MA-LoT achieves 61.07% accuracy, surpassing all baselines, including both tree-search and whole-proof generation methods. These results underscore the advantages of integrating formal verification with structured reasoning leading to better Lean4 theorem proving capability. Beyond theorem proving, LoT-TL training pipeline demonstrates a potential method for applying Long CoT techniques to domain-specific tasks without specialized annotations. Additionally, the success of multi-agent Long CoT in Lean4 suggests broader applications of formal verification for enhancing structured reasoning across diverse fields.